OpenStudy (anonymous):

result of multiplication and addition to eliminate a variable in the system of equations? Of "2x + 3y = 6" and "5x + 2y = 4"

OpenStudy (anonymous):

@terenzreignz:

terenzreignz (terenzreignz):

Of course... Okay, so you need to pick a variable to eliminate... which one do you want to get rid of first? :) (In this case, it doesn't really matter which)

OpenStudy (anonymous):

x?

terenzreignz (terenzreignz):

Okay, so, What do you multiply to 2 so that you get 5?

OpenStudy (anonymous):

You can't multiply 2 by anything to get 5 unless its a decimal.

terenzreignz (terenzreignz):

Okay, let's try another tack... what is the Least Common multiple of 2 and 5 (the coefficients of x in the equations)

OpenStudy (anonymous):

10.......?

terenzreignz (terenzreignz):

Okay, good :) So what do you multiply to 2 to get 10? What do you multiply to 5 to get 10?

OpenStudy (austinl):

We have \[2x+3y=6\] \[5x+2y=4\] We want to get rid of the x, so what I would do is is multiply the top by -5 and the bottom by 2. As such we would have \[-10x-15y=-30\] \[10x+4y=8\]

OpenStudy (anonymous):

you multipy 2 x 5 to get 10 and 5 x 2 to get 10

OpenStudy (austinl):

Sorry, I just kinda jumped, but I already had it written....

OpenStudy (anonymous):

Hahah it's okay.

terenzreignz (terenzreignz):

Okay, so what I want you to do is multiply the first equation (the one with the 2x) with -5 and the second equation by 2....

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

Like use " (,)" ?

OpenStudy (anonymous):

And Distribute?

terenzreignz (terenzreignz):

For instance, multiplying -5 to the first equation, we get: -10x - 15y = -30

OpenStudy (anonymous):

THen for the other equation I would get 10x=4y=8?

terenzreignz (terenzreignz):

YOu mean 10x + 4y = 8, right? :P

OpenStudy (anonymous):

Ohh yeah haha.

terenzreignz (terenzreignz):

Okay, so recap... we now have -10x - 15y = -30 10 x + 4y = 8 Now you can just add these equations together (What this means is form a new equation where the left side is the sum of the left-sides and the right, likewise)

OpenStudy (anonymous):

So I'd have -10x-15y+10x+4y=8-30.....

terenzreignz (terenzreignz):

mhmm :) You'll notice that the x would... cancel out... (get eliminated)

OpenStudy (anonymous):

Oh!

terenzreignz (terenzreignz):

So.. simplify the equation, the x would cancel out, and you can simply solve for y :)

OpenStudy (anonymous):

So far I have -15y+4y=8-30=10-10

OpenStudy (anonymous):

Is that right?

terenzreignz (terenzreignz):

-15y + 4y = 8 - 30 Solve for y.

OpenStudy (anonymous):

Ohh I take out the 10's?

terenzreignz (terenzreignz):

They disappear... they disappear, remember? Because they cancelled out :P

OpenStudy (anonymous):

I'm sorry I'm really stupid I swear i'm not blonde.

terenzreignz (terenzreignz):

:) Just do it :D

OpenStudy (anonymous):

I got y+y= 8-30+15-4..

terenzreignz (terenzreignz):

You don't do that :D You add similar terms...

OpenStudy (anonymous):

What is that....

terenzreignz (terenzreignz):

-15y + 4y = -11y :P This is basic stuff D:

OpenStudy (anonymous):

I'm sorry!!!! D:

terenzreignz (terenzreignz):

so you get -11y = -30+8 Simplify the right side?

OpenStudy (anonymous):

Wait so I subtracted -15 and 4 and kept the 1 y?

terenzreignz (terenzreignz):

Yes :P think of it this way... we factored out the y: \[\large -15y + 4y = y(-15+4)=y\cdot -11 = -11y\]

OpenStudy (anonymous):

OHhh okay well how did I go from subtracting 30 to adding 30?

terenzreignz (terenzreignz):

huh?

OpenStudy (anonymous):

you put -11y= 8+30 when before it was -30 because it was a negative from multiplying -5 and 6.

terenzreignz (terenzreignz):

NO.... I never put up 8+30 it's 8 - 30 :P look again :D

OpenStudy (anonymous):

Ohh sorry haha.

terenzreignz (terenzreignz):

Okay, so we have -11y = 8 - 30 Simplify /solve?

OpenStudy (anonymous):

So I got -11y=-22 then I divided -22 by -11 and got positive 2 so y=2

OpenStudy (anonymous):

Is that right!? :D

terenzreignz (terenzreignz):

yes, y = 2 :) Now replace y with 2 in any of the two equations you started with and solve for x.